# MoH: Multi-Head Attention as Mixture-of-Head Attention

### v1

@@ B.2 Continue-Tuning LLaMA3-8B
Table E: Comparisons between MoH-LLaMA3-8B and LLaMA3-8B-stage1. MoH-LLaMA3-8B outperforms LLaMA3-8B-stage1 by utilizing only 75% of the attention heads.
@@ Appendix C Additional Experiments
Comparison between MoH-LLaMA3-8B and LLaMA3-8B-stage1.    We divide the training process into two stages. Tab. E shows the comparison between MoH-LLaMA3-8B and the model at the end of the first training stage (LLaMA3-8B-stage1). As shown in Tab. E, MoH-LLaMA3-8B quickly recovers the performance of LLaMA3-8B-stage1 within a training budget of 100B tokens. Notably, in English language tasks, MoH-LLaMA3-8B surpasses LLaMA3-8B-stage1 while using only 75% of the attention heads. However, for Chinese language and math tasks, the recovery performance of the MoH model is not as strong as for English. For example, MoH-LLaMA3-8B achieves an accuracy of 64.4 on CMMLU, compared to 66.0 for LLaMA3-8B-stage1. We attribute this to the fact that the model’s Chinese and mathematical capabilities are primarily established during the first training stage. Since the first training stage uses only 300B tokens, significantly less than the 15T tokens in LLaMA3-8B’s pre-training, the model’s abilities in these areas are not fully stable. In the second training stage, after switching to the MoH model, the model experiences more significant forgetting in Chinese and mathematical tasks. Overall, as shown in Tab. E, MoH-LLaMA3-8B achieves an average accuracy of 64.8% across 14 benchmarks, outperforming LLaMA3-8B-stage1 by utilizing only 75% of the attention heads.
Figure A: Additional visualization of the head load distribution in the final MoH layer. For MoH-ViT-B and MoH-DiT-XL/2, we present the head load distributions for the categories “Basketball”, “Bookshop”, “Cart”, “Husky”, and “Jean”. MoH-ViT-B activates 75% of the attention heads. MoH-DiT-XL/2 activates 90% of the attention heads.
@@ Appendix D Additional Qualitative Analysis
Images Generated from the Proposed MoH-DiT-XL/2 Model.    Fig. C shows samples generated by our class-conditional MoH-DiT-XL/2 model. These results demonstrate the ability of MoH-DiT-XL/2 to generate semantically correct content with accurate spatial relationships.
@@ Appendix E Details of Quantitative Evaluations for LLMs
Figure C: Images generated from the proposed MoH-DiT-XL/2 model. We show samples generated from our class-conditional MoH-DiT-XL/2 model trained on ImageNet at 256$\times$256 resolution. MoH-DiT-XL/2 activates 90% of the attention heads.

### v2

@@ B.4 Continue-Tuning LLaMA3-8B
Table G: Comparisons between MoH-LLaMA3-8B and LLaMA3-8B-stage1. MoH-LLaMA3-8B outperforms LLaMA3-8B-stage1 by utilizing only 75% of the attention heads.
@@
Figure A: Additional visualization of the head load distribution in the final MoH layer. MoH-ViT-B activates 75% of the attention heads. MoH-DiT-XL/2 activates 90% of the attention heads.
@@ Appendix C Additional Experiments
Comparison between MoH-LLaMA3-8B and LLaMA3-8B-stage1.    We divide the training process into two stages. Tab. G shows the comparison between MoH-LLaMA3-8B and the model at the end of the first training stage (LLaMA3-8B-stage1). As shown in Tab. G, MoH-LLaMA3-8B quickly recovers the performance of LLaMA3-8B-stage1 within a training budget of 100B tokens. Notably, in English language tasks, MoH-LLaMA3-8B surpasses LLaMA3-8B-stage1 while using only 75% of the attention heads. However, for Chinese language and math tasks, the recovery performance of the MoH model is not as strong as for English. For example, MoH-LLaMA3-8B achieves an accuracy of 64.4% on CMMLU, compared to 66.0% for LLaMA3-8B-stage1. We attribute this to the fact that the model’s Chinese and mathematical capabilities are primarily established during the first training stage. Since the first training stage uses only 300B tokens, significantly less than the 15T tokens in LLaMA3-8B’s pre-training, the model’s abilities in these areas are not fully stable. In the second training stage, after switching to the MoH model, the model experiences more significant forgetting in Chinese and math tasks. Overall, as shown in Tab. G, MoH-LLaMA3-8B achieves an average accuracy of 64.8% across 14 benchmarks, outperforming LLaMA3-8B-stage1 by utilizing only 75% of the attention heads.
Effect of the Activated Head Ratio.    As shown in Tab. H, activating more attention heads generally leads to improved model performance. These results are intuitive, as activating more attention heads equates to utilizing more parameters and performing additional computations on the input.
Table H: Ablation study on the impact of the activated head ratio. All results are from MoH-ViT-S, by using a training budget of 100 epochs.
Activated Heads 50% 55% 60% 65% 70% 75% 80%
Accuracy (%) 78.32 78.38 78.44 78.50 78.42 78.58 78.78
@@ Appendix D Additional Qualitative Analysis
Additional Visualization of the Head Routing Score Distribution.    We provide additional visualization of the head routing score distribution in Fig. C, Fig. D, and Fig. E. As illustrated in Fig. C, Fig. D, and Fig. E, these head routing scores also vary across categories and task types. This dynamic weighting mechanism allows MoH to adjust the importance of each head in response to different task requirements, further enhancing its flexibility and performance. Besides, we find that the routing scores of shared heads change more across categories than those of routing headers. We consider this because routed heads adapt to different categories by adjusting their activation, while shared heads remain activated all the time. Therefore, shared heads primarily rely on changes in routing scores to adapt to different categories.
Images Generated from the Proposed MoH-DiT-XL/2 Model.    Fig. F shows samples generated by our class-conditional MoH-DiT-XL/2 model. These results demonstrate the ability of MoH-DiT-XL/2 to generate semantically correct content with accurate spatial relationships.
@@ Appendix E Details of Quantitative Evaluations for LLMs
Figure C: Additional visualization of the head routing score distribution in MoH-ViT-B. MoH-ViT-B activates 75% of the attention heads.
@@
Figure D: Additional visualization of the head routing score distribution in MoH-DiT-XL/2. MoH-DiT-XL/2 activates 90% of the attention heads.
@@
Figure E: Additional visualization of the head routing score distribution in MoH-LLM-B. MoH-LLM-B activate 75% of the attention heads.
@@
Figure F: Images generated from the proposed MoH-DiT-XL/2 model. We show samples generated from our class-conditional MoH-DiT-XL/2 model trained on ImageNet at 256$\times$256 resolution. MoH-DiT-XL/2 activates 90% of the attention heads.